OpenStudy (albert0898):

What is the solution set for |x - 6| + 4 = 10

OpenStudy (anonymous):

First you have to subtract the 4 from both sides. And then you drop the absolute value sign and solve like a regular equation. You then have to to it to its negative.

OpenStudy (albert0898):

@claritamontano There are two answers for this. x = 12 that I know. What the other value and how do I find it?

OpenStudy (anonymous):

So just set it to its negative. \[\left| x-6 \right| +4=-10\]

OpenStudy (anonymous):

that would be your second equation.

OpenStudy (albert0898):

Then x would equal -8 I put -8, 12 as my solution set on my test and I got that wrong. That's what I'm trying to figure out...

OpenStudy (albert0898):

@claritamontano

OpenStudy (anonymous):

okay so what were your choices?

OpenStudy (albert0898):

1. {0, 12} 2. {-8, 12} What I put 3. {-12, 0} 4. {-12. -8}

OpenStudy (anonymous):

your answer would be 12 not 012.

OpenStudy (albert0898):

What do you mean?

OpenStudy (albert0898):

I put #2 as my answer, almost everyone in my class put #1 and they got it right.

OpenStudy (anonymous):

wait sorry lol answered the wrong question I was looking at my notes and did that problem give me a sec. can I see you steps on how you solved this.?

OpenStudy (debbieg):

Do NOT change the RHS to the negative until you have isolated the absolute value. E.g., do NOT use |x−6|+4=−10 do get the 2nd solution. Isolate the abs value FIRST, then set up both the positive and negative solutions: |x−6|+4=10 |x−6|=−14

OpenStudy (albert0898):

I did it the same way you did it @claritamontano . x - 6 + 4 = 10 x - 2 = 10 x = 12 x - 6 + 4 = -10 x - 2= -10 x = -8

OpenStudy (albert0898):

My teacher taught "drop the bars and negate one equation, then solve"

OpenStudy (anonymous):

yeah thats what my teacher said also.

OpenStudy (albert0898):

I did just that and got the wrong answer... maybe she graded the test papers wrong :?

OpenStudy (debbieg):

Crap, not sure what all that crazyness is. |x-6|+4=10 |x-6|=6 THEN set up the compound equation: x-6=6 OR x-6=-6

OpenStudy (anonymous):

oh i see thanks debbie i learned something also XD

OpenStudy (debbieg):

"drop the bars and negate one equation, then solve" is fine, but only after you isolate the absolute value. Its because of what abs value MEANS. |x|=6 means that x=6 or x=-6 But |x|+a=6 does NOT mean that |x|+a=6 or |x|+a=-6

OpenStudy (albert0898):

I see how you did that Debbie. But I want to know how to get the answer using my teachers method. Drop the bars. Negate the equation. Solve. Can you show me how you do it, step-by-step please.

OpenStudy (debbieg):

Isolate the absolute value expression (get it all by itself on one side of the = sign). Drop the bars. Negate the equation. Solve. What you are doing is getting the point where you have: |stuff|=a where a is some number. Once that is true, then you can say that stuff=a or stuff=-a WHATEVER the stuff. I always say "put blinders on... don't TOUCH what is inside the abs value sign, just isolate it, then split into 2 equations, one = a and the other = -a".

OpenStudy (debbieg):

So you have |x-6|+4=10 |x-6|=6 THEN set up the compound equation: x-6=6 OR x-6=-6 Now, can you solve each of those? Those are your 2 solutions.

OpenStudy (debbieg):

And dont forget the beauty of algebra - you can ALWAYS CHECK YOUR ANSWER. Notice that your solutions, 12 and -8, do not both work.